### v2

# FlexEdit: Marrying Free-Shape Masks to VLLM for Flexible Image Editing

Abstract
Combining Vision Large Language Models (VLLMs) with diffusion models offers a powerful method for executing image editing tasks based on human language instructions. However, language instructions alone often fall short in accurately conveying user requirements, particularly when users want to add, replace elements in specific areas of an image. Luckily, masks can effectively indicate the exact locations or elements to be edited, while they require users to precisely draw the shapes at the desired locations, which is highly user-unfriendly. To address this, we propose FlexEdit, an end-to-end image editing method that leverages both free-shape masks and language instructions for Flexible Editing. Our approach employs a VLLM in comprehending the image content, mask, and user instructions. Additionally, we introduce the Mask Enhance Adapter (MEA) that fuses the embeddings of the VLLM with the image data, ensuring a seamless integration of mask information and model output embeddings. Furthermore, we construct FSMI-Edit, a benchmark specifically tailored for free-shape mask, including 8 types of free-shape mask. Extensive experiments show that our method achieves state-of-the-art (SOTA) performance in LLM-based image editing, and our simple prompting technique stands out in its effectiveness. The code and data can be found at https://anonymous.4open.science/r/FlexEdit
1Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences
2Shenzhen Technology University
3Georgia Institute of Technology
4Carnegie Mellon University
1 Introduction
Image editing has significantly benefited from the exceptional generative capabilities of diffusion models [13, 4, 29, 5, 15, 25, 7, 2, 30, 22, 32]. Trained on large scale datasets of text-image pairs, these models can generate high-quality images that align with text prompts. Instruction-based image editing extends these capabilities by using human language instructions to modify specific elements within an original image. However, researchers have found that language instructions can often be ambiguous and complex [10, 23, 9], thus the text encoders [24] fail to accurately represent this information, resulting in subpar editing outcomes. Large Language Models (LLMs) [18, 21], with their strong instruction-following capabilities, can understand complex human instructions and reason out clear and simple editing directives. Models like SmartEdit [9] have pioneered the use of LLMs as text encoders to comprehend complex instructions, allowing users to input complex commands for image editing.
Figure 1: Free-shape mask can simplify the user input when dealing with location related instruction, and more flexible for user input. The compared methods are SmartEdit and Largen respectively.
Figure 2: FlexEdit support both (a) single image and (b) multiple image editing, including replace/add an object. The free-shape mask input indicate the localization information, highly user-friendly than asking user to input full mask.
However, when it comes to scenarios that are difficult to describe with language, such as specifying precise locations within an image, user have no idea how to prompt the models. Even if the model can understand, such prompts are clearly not user-friendly (e.g., “$3^{th}$ row and $3^{th}$ column of the grid”). Users spend excessive time crafting these prompts, and the model may still fail to grasp their meaning, leading to unsatisfactory results (see Upper Figure 1). To address these challenges, visual prompts, such as masks, can effectively and clearly indicate the exact locations or elements to be edited, overcoming the limitations of language commands. The challenge of reducing the "tedious" and "time-consuming" effort of manual segmentation is well-documented in HCI literature, with foundational works like Intelligent Scissors [19] and GrabCut [27] created to simplify the process. Our work continues this user-centric philosophy. However, most contemporary methods that utilize mask references still require users to precisely draw shapes at the desired locations, which is highly user-unfriendly and can be a barrier to effective use [4]. If the mask shapes are irregular or do not completely overlap with the objects, they often produce very poor results (see Lower Figure 1).
In real-world scenarios, the shapes users draw for masks are often random and may cover only a small part of the intended area. These free-shape masks challenge the model’s ability to understand user requirements. To address this issue, we created the Free Shape Mask Instruction Edit (FSIM-Edit) benchmark, which includes 80 and 125 single images and multiple images respectively, 8 types of free-shape mask. Evaluating existing models with this benchmark revealed that they perform poorly in free-shape mask setting, even after the fine-tuning. To overcome these issues, we modified the mask functionality in MagicBrush [32], integrating complex instructions with image editingin the FSIM-Edit dataset and proposed FlexEdit, an innovative end-to-end image editing method that leverages both free-shape masks and language instructions for Flexible Editing (FlexEdit). FlexEdit employs a Vision Large Language Model (VLLM) to comprehend the image content, visual prompts, and user instructions. We introduced an adapter structure that fuses the mask represented embeddings of the VLLM with the image data, ensuring a seamless integration of visual information and model output embeddings. This integration enhances the model’s ability to understand and execute complex editing tasks accurately. Experiments demonstrate that FlexEdit significantly outperforms other methods under free-shape mask conditions. Notably, our simple prompting technique proves to be highly effective, setting a new standard for ease of use and precision in the field. FlexEdit represents a significant step forward in the realm of intelligent image editing, combining the strengths of VLLMs and diffusion models to provide a user-friendly and highly capable editing solution. Our contributions can be summarized as follows:
  • We propose FlexEdit, an end-to-end image editing method that combines free-shape masks and language instructions, overcoming the limitations of traditional methods that require precise mask drawing.
  • We introduce the Mask Enhanced Adapter (MEA) structure that seamlessly enhance the mask embedding of the Vision Large Language Model (VLLM) with image data, enhancing the model’s ability to understand and execute complex editing tasks.
  • We create the Free Shape Mask Instruction Edit (FSIM-Edit) benchmark, which includes a comprehensive dataset with diverse scenarios and editing instructions, to rigorously evaluate the performance of image editing models under free-shape mask conditions.
2 Related Works
In this section, we review the related works in image editing using diffusion models and the advancements in powerful Vision Large Language Models in image editing.
2.1 Image Editing with Diffusion Models
While our work is situated within the paradigm of VLLM-guided diffusion models, it is worth noting the contributions of Generative Adversarial Networks (GANs) to image editing. GAN-based methods often offer faster inference speeds. However, diffusion models generally provide superior training stability, mode coverage, and higher generation fidelity, which are critical for the complex, instruction-guided editing tasks that we target. Therefore, we focus on diffusion models and their direct state-of-the-art competitors for the most relevant and fair comparison.
Image editing has achieved exceptional performance thanks to the generative abilities of diffusion models [2, 32, 7, 4, 22, 30]. [2] introduced InstructPix2Pix, an instruction-based image editing approach that modifies images based on human language instructions. To better align with human instructions, [32] proposed MagicBrush, extending text-only instructions to scenarios where a visual mask is provided. On the other hand, subject-driven image editing tasks aim to generate images conditioned on a customized subject and a text prompt that describes the context. For example, Anydoor [4] can teleport target objects to new scenes at user-specified locations (mask) with desired shapes. [22] proposed Largen for image inpainting that enables seamless inpainting of masked scene images, incorporating both the textual prompts and specified subjects. [28] and [29] start discussing to use coarse visual prompt such as bounding boxes, scribble in guiding the editing, but they use the classic text encoder in diffusion models, which makes it hard to understand those complex instructions. These models can combine user-provided images with visual prompts, such as masks, to perform image editing. The masks they use need to precisely overlap with the area to be edited, which is not user-friendly. In comparison, we propose using a free-shape mask as a visual prompt and introduce vision large language models (VLLMs) to better understand complex scenarios and instructions for visual prompt-based image editing in an end to end manner.
2.2 VLLMs in Image Editing
Vision Large Language Models (VLLMs) are known for their extensive world knowledge and ability to understand complex instructions [1, 14, 17, 16, 11, 3]. Fine-tuning these VLLMs for image generation tasks has shown great success in comprehending human language instructions and generating images [9, 6, 13]. For example, [9] fine-tune the LLaVA [17] model with dataset of MagicBrush [32] to fit it into the image editing domain. While InstructAny2Pix [13] build a multi-modal editing system that enables users to edit an image using instructions involving audio, images, and text. Despite their success, current models predominantly rely on language-based instructions, overlooking the power of visual prompts such as masks. This reliance can limit the intuitiveness and accessibility of the editing process. Our approach seeks to enhance the editing capabilities by incorporating both user instructions and visual prompts, thereby achieving more precise and user-friendly image edits.
Figure 3: The architecture of the FlexEdit framework. FlexEdit integrates visual prompts and human instructions for complex image editing. It utilizes a VLLM backbone for multi-modal instruction understanding, a Q-Former for refining the hidden states, and a Mask Enhanced Adapter (MEA) for merging image and language model outputs. The final image generation is achieved through a diffusion model.
3 Methodology
This paper presents FlexEdit designed to integrate human instructions and visual prompts for complex image editing tasks. Existing methods, whether based on language instructions or subject mask editing, struggle with understanding user-provided free-shape masks [4, 9]. This challenge lead to a situation that user struggle to draw a mask to meet the models’ need, which is not user-friendly and not acceptable.
In this section, We begin by introducing the FlexEdit Framework and explaining how it integrates free-shape mask with human instructions. We then discuss the Mask Enhanced Adapter (MEA) Module, which plays a crucial role in fusing image features. Finally, we explain the training process for understanding free-shape masks and present the FSMI-Edit Benchmark for evaluating our method.
3.1 FlexEdit Framework
The framework processes three inputs: a scene image $x_{1}$ with an associated mask $m$, a subject image $x_{2}$, and an editing instruction $\tau$. The primary objective is to generate a target image that incorporates $x_{2}$ and $\tau$ as the content, while utilizing $m$ as a positional reference within $x_{1}$. Importantly, the mask $m$ can take on a free-form shape and does not need to fully overlap with the intended position.
We expanded the vocabulary of LLaVA [17] with $N$ new tokens “<$img0$>, …, <$imgN$>” to represent image editing information. As shown in Figure 3, $x_{1}$, $x_{2}$, $m$, and $\tau$ are input into the Vision Large Language Model (VLLM) $V(\cdot;\omega)$ to obtain response tokens $r_{i}$. Following this, the response set $R=\{r_{1},r_{2},\dots,r_{n}\}$ is processed to extract the corresponding hidden states $h_{i}\in\mathcal{H}$, capturing the deeper semantic information from the VLLM’s understanding of the inputs. However, these hidden states primarily reside in the LLM’s text vector space, which poses compatibility challenges when interfacing with the diffusion model-particularly one trained on CLIP [24] text embeddings. To address this, we introduce a Q-Former [12] module that refines the hidden states into embeddings $e$ compatible with the diffusion model. The transformation process is summarized as follows:
$R{}$ $=V(x_{1},x_{2},\tau,m;\omega)$ (1)
[…]
$h_{i}$ $=H(x_{1},x_{2},\tau,m;\omega|R^{t}_{i})$
$e{}$ $=Q(\mathcal{H})$
where $R$ is the response set, $R^{t}$ is the set of image tokens, and $e$ represents the embeddings transformed by the Q-Former function $Q$.
For efficient training, we utilize LoRA [8] fine-tuning techniques, freezing most VLLM parameters. Let the ground truth text label of the VLLM response be $R^{\prime}$. The VLLM optimization process follows the recurrent loss calculation:
$\mathcal{L}_{\text{VLLM}}(x,\tau)=-\sum_{\{(x,\tau),R^{\prime}\}}\log p_{\omega+\Delta\omega(\theta)}(R^{\prime}\mid(x,\tau))$ (2)
where $\theta$ represents the LoRA parameters.
Subsequently the Mask Enhanced Adapter (MEA) $A(\cdot;\delta)$ takes embedding $e$ and images $x_{1}$, $x_{2}$ as input, enhancing the mask feature from the VLLM with scene and subject images. The output $c$ from the MEA UNet block serves as the key and value for cross-attention in the subsequent block recurrently, calculated as:
$c=A(x_{1},x_{2},e;\delta)$ (3)
here $A(\cdot;\delta)$ is the MEA function with parameters $\delta$, and $c$ is the MEA UNet block’s output.
During the diffusion process, the encoded image latent $z=\mathcal{E}(x_{1})$ is concatenated with the noisy latent $z_{t}$ and fed into the UNet [26] in a residual manner. The UNet $\epsilon_{\alpha}$ is trained to predict the noise added to the noisy latent $z_{t}$, with noise levels increasing under timesteps $t$, given the condition $c$ from the MEA module. The diffusion optimization process is formulated as:
$\begin{split}\mathcal{L}_{\text{diffusion}}&=\mathbb{E}_{\mathcal{E}(y),\mathcal{E}(x_{1}),c_{T},\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\\ &\epsilon_{\alpha}(t,\text{concat}[z_{t},\mathcal{E}(x_{1})],c)\|_{2}^{2}\Big{]}\end{split}$ (4)
where $\epsilon$ is the unscaled noise, $t$ is the sampling step, $z_{t}$ is the latent noise at step $t$, $\mathcal{E}(x_{1})$ is the encoded scene image, and $c$ is the output from the MEA module.
The overall loss function combines VLLM and diffusion losses:
$\mathcal{L}_{\text{overall}}=\mathcal{L}_{\text{diffusion}}+\mathcal{L}_{\text{VLLM}}$ (5)
3.2 Mask Enhanced Adapter
The Mask Enhanced Adapter (MEA) Module aims at fusing the mask editing information from VLLM with the scene and subject images, enhancing the feature for fully interaction. Passing them to the diffusion models allows it to generate the result image for users.
The scene $x_{1}$ and subject $x_{2}$ images are first mapped to the text embedding space via the CLIP image encoder, for CLIP’s training objective is to maximize the similarity of the image and text label embedding. The output $e$ of QFormer contains the editing intent generated by VLLM after combining the mask and instruction. However, it lacks fine-grained details from the original image. To reintroduce this information, we employed a cross-attention mechanism to fuse the output $e$ from QFormer [12] with the features from the scene image $x_{1}$, resulting the scene features $f_{1}$. The key and value of the cross-attention come from the scene image $x_{1}$ with the the output $e$ from QFormer as the query. Formulated as follow:
$f_{1}=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V$ (6)
where $Q=eW_{q1}$, $K=x_{1}W_{k1}$, $V=x_{1}W_{v1}$ are the query, key, and values matrices of the attention operation, and $W_{q1}$ , $W_{k1}$, $W_{v1}$ are the weight of the trainable linear projection layers.
Considering that the the edited part in final image need to closely resemble a specific object in subject image $x_{2}$, we utilized the linear layer and the decouple cross attention to integrate these features separately, rather than directly fusing them through cross-attention. The subject feature is obtained by mapping the text embedding from $x_{2}$ through a linear layer. And then the features $f_{1}$ and $f_{2}$ go into the UNet block with decouple cross attention where the cross-attention layers. We add a new cross-attention layer for each cross-attention layer in the UNet block to for features interaction. Finally, given the query features $Z$, scene features $f_{1}$ and subject features $f_{2}$, the output of cross-attention $Z^{\prime}$ can be defined by the following equation:
$Z^{\prime}=\text{Softmax}\left(\frac{QK_{1}^{\top}}{\sqrt{d}}\right)V_{1}+\lambda\text{Softmax}\left(\frac{QK_{2}^{\top}}{\sqrt{d}}\right)V_{2}$ (7)
where $Q=ZW_{q2}$, $K_{1}=f_{1}W_{k2}$, $V_{1}=f_{1}W_{v2}$ and $K_{2}=f_{2}W_{k2}$, $V_{2}=f_{2}W_{v2}$ are the query, key, and values matrices of the attention operation respectively, and $W_{q2}$ , $W_{k2}$, $W_{v2}$ are the weight matrices of the trainable linear projection layers. $\lambda$ is weight factor, and we set $\lambda=0$ when editing single image.
3.3 Training to Understand Free-Shape Mask
For easy annotation and efficient training, we modified the classical segmentation dataset COCO [31] and GRefCOCO [15] for free-shape mask understanding training process. Specifically, we first adapt the original mask label $m_{o}$ into free-shape mask $m=W(m_{o})$ by a simple random walking algorithm $W$, and model are trained to predict the original mask while output the label of the predict class $p$. For example, the output can be “The masks means ‘chair’, <$img0\dots N$>”, <$img0\dots N$> represent the generate image with original mask $m_{o}$ (see Figure 4). More technical details will be discussed in the supplementary material.
Figure 4: We train model to understand free-shape mask by having it predict the full mask from a given free-form mask. The picture shows two examples of this process.
3.4 FSMI-Edit Benchmark
To address the lack of standardized evaluation for this task, we construct the Free-Shape Mask Instruction Edit (FSMI-Edit) benchmark. To our knowledge, it is the first benchmark specifically designed to evaluate image editing under the realistic condition of imprecise, free-shape masks. We focus on daily indoor (e.g. Backpack, Kitchen) and outdoor (e.g. Animals, Ground) scenarios, for subject image we also consider two types of both simple (only subject provided) and complex (subject with noisy background) subject image. FSMI-Edit consists of 80 and 125 single images and multiple images editing picture respectively. We carefully design the mask type so that it can mimic human input preference, thoroughly evaluating the performance of free-shape mask reference image editing. Example of the free-shape mask from FSMI-Edit is shown at Figure 5. The free-shape mask consists of 8 type total: circle, circle(open hole), rectangle, rectangle (open hole), triangle, triangle (open hole), irregular, irregular (open hole) mask, with different orientation. We hope more researchers will pay attention to the free-shape mask guided image editing from these perspectives, thereby fostering the user-friendly multi-modal instruction image editing methods.
Figure 5: Example of the different types of the free-shape mask from FSMI-Edit benchmark.
4 Experiments
Table 1: Comparison in multiple image editing domain, we compare the results in full mask and free-shape mask setting.
Method Foreground Whole Image PNSR $\uparrow$ CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINOv2 $\uparrow$ PNSR $\uparrow$ LPIPS $\downarrow$ CLIP-I $\uparrow$ DINOv2 $\uparrow$ Full Mask AnyDoor 13.1316 22.0478 0.8959 0.2485 19.2274 0.1055 0.2340 0.9292 Largen 12.5130 20.8105 0.8332 0.2468 18.6747 0.1166 0.2331 0.8886 InstructAny2Pix 8.6129 20.9438 0.7557 0.2424 10.3434 0.5231 0.1723 0.7100 FlexEdit 13.1645 24.1754 0.8279 0.2982 18.0522 0.1433 0.2402 0.8596 Free-Shape Mask AnyDoor 11.8483 19.5245 0.7711 0.2451 18.4441 0.1348 0.2280 0.8292 Largen 12.1133 19.1737 0.7627 0.2386 18.6563 0.1373 0.2266 0.8271 InstructAny2Pix 9.6056 20.5750 0.7582 0.2411 11.4099 0.4720 0.1775 0.7151 FlexEdit 12.5648 22.4813 0.7795 0.2908 17.9068 0.1585 0.2353 0.8262
We compare our FlexEdit with SOTA models in both the single and multiple image editing domains. For a fair comparison, we fine-tune the models on the same dataset we use. For models that cannot be fine-tuned, we also report their performance under standard mask settings to assess their ability to understand the content when mask provided.
4.1 Experiments Setting
4.1.1 Training Datasets
Figure 6: Qualitative results for free-shape mask guided image editing in multiple image setting.
We include training data from four categories, with a large-scale, mixed composition: (1) 65To address the imprecision of user-drawn masks in real-world applications, we avoid using the actual masks directly as inputs. Instead, we simulate user input by generating free shape masks using random walks on the real masks. These generated masks are then reviewed and re-annotated by three volunteers to ensure they closely mimic real-life user inputs. For the segmentation datasets, the model’s objective is to regenerate masks based on the free-form input to match the ground truth shapes. Training on the VQA dataset ensures that the model retains its analytical and reasoning capabilities during fine-tuning. Additional technical details can be found in the supplementary materials.
4.1.2 Metrics
The final objective for the model is to focus understand the free-shape mask area during editing, adhering to the instructions while minimizing changes to the overall image. To assess the ability in the mask understanding, we modified the mask-evaluate setting. Unlike the setting of aforementioned methods that use the mask area in evaluating the generate quality, given that free-shape masks’ imprecise nature, we use the real edited regions from the GT as the evaluate masks instead of the input one. In thoroughly assessing the quality of image editing, we split the metric into the single and multiple image evaluation:
Multiple Image: We include PNSR, CLIP-T, CLIP-I [24], and DINOv2 [20] for the metric of foreground quality assessing, where CLIP-T measures how performance of the edited foreground aligns with the text label. PNSR, CLIP-I and DINOv2 assess the visual similarity between the generated images and the GT foreground images. LPIPS was introduced in the whole image evaluation while CLIP-T is excluded since CLIP-T label can not be annotated in the whole image setting.
Single Image: For single-image editing, in line with SmartEdit [9], we use PSNR, SSIM, LPIPS to evaluate background similarity, and CLIP-T to assess the alignment of the results’ foreground with the text label.
4.2 Comparisons with SOTA models
4.2.1 Baselines
We compared the diffusion only method AnyDoor and Largen, as well as the VLLM plus diffusion method instructAny2Pix, with our method in multiple image editing. Here, AnyDoor and Largen are the methods designed for zero-shot subject-driven image editing task, generating appropriate objects from the subject image within the masked area of the scene image. InstructAny2Pix leverages a VLLM to combine multiple user images and text instructions for editing the corresponding images.
For single image editing, we compare FlexEdit with InstructPix2Pix, InstructAny2Pix, and SmartEdit. InstructPix2Pix and SmartEdit edit images based on human language instructions. InstructPix2Pix uses a CLIP structure to understand human instructions and image content, while SmartEdit employs a large visual language model to interpret human instructions and image content.
4.2.2 Multiple Image Editing
We first provide a qualitative comparison between FlexEdit with other method with add/replace operation and simple/complex subject images in Figure 6. The compared model either fail to understand the free-shape mask (AnyDoor, Largen), or unintentionally change the background image (InstructAny2Pix).
Foreground Quality Comparison: Under the normal mask setting, the AnyDoor and Largen shows the best score in CLIP-I, indicating the exhibit of the highest similarity to the foreground GT image, wheras InstructAny2Pix is less than ours. DINOv2 shows the same trend as CLIP-I, except ours FlexEdit achieved 0.2982, marginly higher than others. For the CLIP-T scores which represented the similarity between the generated foreground and the text label, however, FlexEdit achieved 24.1754, achieving the SOTA results. In the free mask setting, where user-provided masks are more arbitrary and abstract, methods like AnyDoor and Largen, which generate images only within the mask, fail to meet user expectations, resulting in a significant drop in metrics. InstructAny2Pix’s performance do not decrease and perform stable in this setting. In this setting, our model have a sightly performance drop, while achieve the state-of-the-art results among all the metrics, showing our strong ability in understanding the free-shape mask. The detail results can be found at Table 1 (Foreground Image).
Whole Image Quality Comparison: The PNSR, CLIP-I, DINOv2 is close in the whole image setting for AnyDoor, Largen and FlexEdit. Note that InstructAny2Pix perform marginly bad at this setting, indicating that it will modify the unintentional background image, this is a long-stand problem in an end-to-end diffusion generation framework. However, our FlexEdit while edit the foreground object well, keeping the whole image stable, demonstrating our model’s strong image understanding and image edting abilities. More detail results can be found at Table 1 (Whole Image).
4.2.3 Single Image Editing
For a fair comparison, we fine-tune InstructPix2Pix and SmartEdit under the same setting as ours [^1] InstructAny2Pix has not released the training code at the time this paper submitted.. The qualitative comparison is shown in Figure 7, InstructAny2Pix can understand the free-shape mask, while will change the background image unintentionally. SmartEdit works better than InstructPix2Pix for it can understand the mask occasionally, instructionPix2Pix however, shows bad performance among the three methods.
Figure 7: Qualitative results comparison on single image edit.
Results shows that InstructPix2Pix performs bad even after fine-tuning, we argue that these can come from the problem that CLIP simple encoder can not fully understand the free-shape mask. InstrucAny2Pix still suffers from it can not maintain the whole image during the editing process. FlexEdit in the other hand, understand the free-shape mask and keep the rest of the image still, this can come from our MEA module for its ability in information extraction. SmartEdit shares the similar result as ours, with a sightly weakness that can come from our single-multiple image mixed training strategy. Detail results can be found at Table 2. More comparison can be found in the supplementary material.
Table 2: Comparison in single image editing domain.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CLIP-T $\uparrow$ InstructPix2Pix 19.9414 0.8016 0.1043 20.7010 InstructAny2Pix 15.5149 0.5919 0.2751 20.3758 SmartEdit 21.9535 0.7620 0.0781 20.4910 FlexEdit 22.8081 0.7712 0.0727 20.7081
Table 3: Ablation study in multiple image editing. We evaluate the ‘CA’, standing for cross attention in MEA, and ‘DCA’, standing for decouple cross attention.
CA DCA Foreground Whole Image PNSR $\uparrow$ CLIP-T $\uparrow$ CLIP-I $\uparrow$ DINOv2 $\uparrow$ PNSR $\uparrow$ LPIPS $\downarrow$ CLIP-I $\uparrow$ DINOv2 $\uparrow$ 12.5794 22.4391 0.7796 0.2890 17.9068 0.1585 0.2353 0.8262 - 12.1836 21.6230 0.7451 0.2846 17.5058 0.1806 0.2378 0.8140 - 11.8367 21.3692 0.7561 0.2799 17.0976 0.1813 0.2341 0.8124
4.3 Ablation study on MEA
To validate our contribution, the ablation study was intentionally focused on the MEA’s key components: the Cross Attention (CA) and Decouple Cross Attention (DCA) modules. This follows the standard experimental practice of isolating variables to clearly demonstrate the effectiveness of a new component. We designed two variant structures and compared their performance metrics in both single and multiple image editing to validate the effectiveness of each module. First, we remove the Cross Attention (CA) layer and directly used the output of the QFormer as the scene feature input into the following decouple cross attention, without the scene picture interaction. We also change the Decouple Cross Attention (DCA) structure with a general cross attention and examine the MEA effectiveness in image fusion. As shown in Table 3, in multiple image editing, the CA and DCA can help boost the model performance, both seperately and jointly. For single image editing, as shown in Table 4, though the MEA module does not work in this mode, it can learn during the joint training process, and CA, DCA involving shows a better result for the single image performance.
Table 4: Ablation study in single image editing, same setting as experiment in multiple image editing.
CA DCA PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ CLIP-T $\uparrow$ 22.8081 0.7712 0.0727 20.7081 - 21.8579 0.7546 0.0931 19.8630 - 20.4195 0.7230 0.1186 19.8566
4.4 Limitations and Failure Case Analysis
Despite the strong performance of FlexEdit, our work has some limitations and potential areas for future improvement.
Choice of VLLM Backbone. We chose LLaVA to ensure a controlled and fair comparison with SOTA methods like SmartEdit, allowing us to isolate the benefits of our proposed contributions. While this provides a clear comparison, future work could explore performance improvements with alternative or more recent VLLM backbones, such as Qwen-VL[1].
Failure Cases. We have identified three primary modes of failure:
  • VLLM Grounding Failure: When a free-shape mask is highly ambiguous (e.g., covering two distinct objects simultaneously), the VLLM may misinterpret the user’s intent, leading to an incorrect edit.
  • Semantic Misinterpretation: In some cases, the model may generate an object that is semantically related to the instruction but is contextually incorrect (e.g., generating a tiger when a cat was requested).
  • Background Artifacts: Minor, unintended changes to the background style or texture can occasionally occur. This is a known challenge in end-to-end diffusion generation frameworks that can affect the overall image quality.
5 Conclusion
In this paper, we introduce FlexEdit, an end-to-end method based on Visual Language Models (VLLM) for free shape mask guided image editing. Our model addresses the challenge of imprecise and irregular mask shapes drawn by humans by incorporating free shape masks during the training process. To achieve effective image fusion, we propose a Mask Enhanced Adapter (MEA), which enhanced the mask embedding from the VLLM with image data for seamless blending. To support the image editing community, we also develop FSMI-Edit, a new evaluation dataset specifically designed for free shape mask scenarios. Our method achieves state-of-the-art results on the FSMI-Edit benchmark, outperforming other models trained under similar conditions, whether in free shape mask or full-mask settings.